OpenStudy (anonymous):

Please help!! Will fan and medal!!! Rewrite the following quadratic function in vertex form. Then, determine the axis of symmetry. y = 5x^2 + 15x - 2

OpenStudy (anonymous):

Do you know what vertex form is?

OpenStudy (anonymous):

Not really. I know that it is y = a(x – h)^2 + k

OpenStudy (anonymous):

f (x) = a(x - h)2 + k, where (h, k) is the vertex of the parabola.

OpenStudy (anonymous):

Yeah. but idk how to put it into that form

OpenStudy (anonymous):

Your equation is y=5x^2 + 15x - 2. The first thing you do is add 2 to both sides so it cancels out on the right, making y+2 = 5x^2 + 15x.

OpenStudy (anonymous):

Next, you factor out the 5x^2. This looks like : 5(x^2 + 5x).

OpenStudy (anonymous):

This is called "Completing the square".

OpenStudy (anonymous):

Now, we have to deal with an additional variable, "y" ... so we cannot "get rid of " the factored 5. When we add a box to both sides, the box will be multiplied by 5 on both sides of the equal sign.

OpenStudy (anonymous):

The equation now looks like : y+2(5) = 5(x^2 + 5x + ( 5)

OpenStudy (anonymous):

Still with me @madison.bush